A trapezoid is formed by lines with the following equations
2x+4y=16 x=4 x=0 y=0
Find the area of the trapezoid.

Answer by longjonsilver(2297)

area of trapezoid is (h/2)(a+b) where h is the perpendicular distance between the 2 parallel sides, which have lengths a and b.

2 parallel lengths are those defined by x=0 and x=4... how long are they though? well they intersect the line 2x+4y=16, so...

when x=0, 2x+4y=16 gives 2(0)+4y=16
0 + 4y = 16
4y = 16
y = 16/4
y = 4

when x=4, 2x+4y=16 gives 2(4)+4y=16
8 + 4y = 16
4y = 8
y = 8/4
y = 2

So, the 2 lengths for a and b are 4 and 2

The perpendicular length between them is 4... the distance between x=0 and x=4

So, put these values into the formula above.

jon.